the curve y = f(x) has a y-intercept of 3. the tangent to this curve at the point (x, y) has an x-intercept of x+2. What is the equation of this curve?


1
Expert's answer
2022-03-07T10:07:01-0500

The slope of the tangent to the curve at the point "(x, y)"  is


"y'=x+2"

Integrate


"y=\dfrac{1}{2}x^2+2x+C"

The curve "y = f(x)" has a y-intercept of 3


"3=0+0+C"

"C=3"

The equation of this curve is

"y=f(x)=\dfrac{1}{2}x^2+2x+3"